OpenStudy (racelynn):

how do you work out the problem: 2m+0.5(m-4)+9

OpenStudy (anonymous):

multiply 0.5 to (m -4) then solve for m

OpenStudy (anonymous):

2m+0.5(m-4)+9 2m+0.5m-2+9 then combine like terms and there you have it

OpenStudy (racelynn):

then what

OpenStudy (anonymous):

what do you mean then what there is no equal sign

OpenStudy (racelynn):

i meant it equals 9 not plus 9 sorry

OpenStudy (anonymous):

okay here 2m+0.5m-2=9 2.5m-2=9 +2 +2 2.5m=11 then divide 11 by 2.5 and get 4.4

OpenStudy (anonymous):

so m=4.4

OpenStudy (racelynn):

thanks that makes more sense to me now

OpenStudy (anonymous):

your welcome